What volume of 0.0100 M NaOH must be added to 1.00 L of 0.0500 M HOCl to achieve a pH of 8.00? Ka for HOCl is 3.5 ́ 10–8.
1
Expert's answer
2019-11-18T06:27:58-0500

Solution.

"HClO + OH^- = H2O + ClO^-"

"n(HClO(init)) = 0.05 \times 1.0 = 0.05 \ mol"

"n(NaOH(init)) = 0 \ mol"

"n(ClO^-(init)) = 0 \ mol"

Once we have added a certain amount of alkali, the amounts of substances will be as follows:

"n(HClO(fin)) = 0.05 - x \ mol"

"n(NaOH(fin)) = x \ mol"

"n(ClO^-(fin)) = x \ mol"

Let us use the Henderson-hasselbach equation:

"pH = pKa + lg(\frac{[ClO^-]}{[HClO]})"

"pKa = -lg(Ka) = 7.46"

"8.00 = 7.46 + lg(\frac{x}{0.05-x})"

"3.47 = \frac{x}{0.05-x}"

x = 0.039 mol

"C = \frac{n}{V}"

"V = \frac{n}{C}"

"V = \frac{0.039}{0.01} = 3.9 \ L"

Answer:

V(NaOH) = 3.9 L